Help on this question of Linear Algebra, thanks.

This is a continuation of the previous problem. Compute lim A PO lim A4) k+00 where ро 0.8 0.2

A = 0.7 0.1 0.3 0.9 :)
Answer #1

A 0.7 2:3 2.9 jo [O -1] [39 to [an] - to A= A lok Ak. |A, -A) = 8-) (9-2)-3 17-1 9-) ♡ 9 o a 2-167tboco 7 7 2²-100 -62 tbocoSo Da 6 for, are A x = - ax 39 2 7 Tout = 64. of my tego Bret qua 6m 3 ay t3m to se try o » Charm (my, ay) (i (rou, -ry) (1,family- Genus Species- on Twith = 10m 7-364+Myco 3 mt q2 = 10M2 or 3 m - m 0 mm = 34. Comics) = (–1734) SoIP= (13) cr, ca toDate: AK lok [142] 23 21:21 9 64 elok Exotick 3.6 ch Grook 34104 of 3.6 ch 3x10h 46k richotos 10k + 3.6k 3 +(4.) altek - 3+383 Colin kych 4 72 -3 Ak ) 4 ) Clim 0.8 Pos 3 Kyeb 9 ra 3 1 9 2.9.0.2 -2.4+02 1 2.2 4 2.2. eiss ClimAh) po -0.55 CONTECAS C.Sc